OpenStudy (anonymous):

How do you solve 6(5m-3) = 1/3(24m+12) ?

OpenStudy (anonymous):

are you solving for m?

OpenStudy (anonymous):

1) distributive property 2) combine like terms 3) Solve :)

OpenStudy (anonymous):

Distribute 30m-18=8m+4, then subtract 8m to both sides 22m-18=4; 22m=4+18; 22m=22; m=1

OpenStudy (anonymous):

Thank You!

OpenStudy (anonymous):

can any of you guys help me with a problem?